# Planning and Reasoning with 3D Deformable Objects for Hierarchical Text-to-3D Robotic Shaping

### v1

@@ IV-C Human Survey
TABLE II: Full Pipeline Evaluations. Image and text CLIP score is between 0 and 1, while the human evaluations are between 1 and 5.
@@
CLIP Human Eval. Prompt Human Eval. Quality CLIP Human Eval. Prompt Human Eval. Quality
X 0.278 $\pm$ 0.014 3.91 $\pm$ 0.68 3.57 $\pm$ 0.73 0.295 $\pm$ 0.008 4.67 $\pm$ 0.47 4.52 $\pm$ 0.59
line 0.271 $\pm$ 0.009 2.76 $\pm$ 0.97 2.43 $\pm$ 0.73 0.289 $\pm$ 0.002 4.00 $\pm$ 0.69 3.81 $\pm$ 0.79
flower 0.294 $\pm$ 0.004 2.33 $\pm$ 0.94 2.10 $\pm$ 0.97 0.302 $\pm$ 0.021 3.57 $\pm$ 1.05 3.81 $\pm$ 1.00
column 0.270 $\pm$ 0.003 2.62 $\pm$ 1.00 2.52 $\pm$ 0.91 0.306 $\pm$ 0.005 4.00 $\pm$ 0.93 3.86 $\pm$ 0.94
pyramid 0.275 $\pm$ 0.009 1.91 $\pm$ 0.87 1.76 $\pm$ 0.68 0.289 $\pm$ 0.007 3.67 $\pm$ 0.94 3.48 $\pm$ 1.05
all shapes 0.278 $\pm$ 0.012 2.71 $\pm$ 1.12 2.48 $\pm$ 1.02 0.296 $\pm$ 0.013 3.98 $\pm$ 0.93 3.81 $\pm$ 0.98
Figure 5: The human oracle is required to follow the same process of placing discrete chunks of clay in the scene, before slowly sculpting the desired shape. We chose to visualize the results as point clouds to get the best visualization of the full 3D geometry.
@@ IV-D Text-to-3D Shape Goals
To fully evaluate our proposed system, we conducted three real-world experiments for each shape prompt for the ’X’, ’line’, ’flower’, ’column’, and ’pyramid’ shape goals. A visualization of the final shapes created by our system is shown in Figure 5. As we present the first text-to-3D real-world shaping system, there are no existing baselines to compare to, thus we focus on comparisons to a human oracle. The quantitative results of the final CLIP cosine similarity and human evaluations for our system compared to a human oracle are shown in Table II. The shapes created by our pipeline are often visually similar to that of the human oracle, but with some imperfections such as variable z-height on the surface caused from the ridges created by the robot’s grasps, etc. We can see from the image-based CLIP scores, the human oracle creates higher quality shapes than our proposed method (Welch’s t-test, $p=0.0006$, $t=3.86$), though in absolute CLIP value our proposed method performs comparably for the ’X’, ’line’, and ’flower’ shapes. These patterns hold for human evaluation scores as well, with the human oracle being rated higher in both prompt similarity (Welch’s t-test, $p=2.33e^{-16}$, $t=8.95$) and shape quality (Welch’s t-test, $p=1.91e^{-18}$, $t=9.65$). While there is a clear statistical significance between the human oracle and full pipeline performance on both human rating metrics, the human oracle itself did not receive a near-perfect score. The survey respondents struggled to quantify the shapes ’pyramid’ and ’flower’. It is very difficult to quantify the concept of how much a shape is a particular shape. To better investigate this difficulty, we conducted a Krippendorff analysis [42] of the inter-annotator agreement across all shapes created by both the human oracle and our pipeline. For the ordinal data of our individual Likert-type questions, the alpha scores for the shape quality and prompt similarity were $\alpha_{quality}=0.476$, and $\alpha_{prompt}=0.460$, respectively. For the Krippendorff analysis, a score of 0 signifies that the level of agreement is consistent with randomly answering the questions, while a score of 1 means the reviewers perfectly agreed. The scores for both of our questions were approximately 0.5, which demonstrates that the responses are not random, but the ratings have substantial noise and the reviewers are not able to fully agree. This further emphasizes the necessity to explore better strategies to quantify performance for the clay shaping task in future work, as even human evaluations have disagreement.
@@ IV-E Semantically Tuning Shape Goals
Figure 6: By semantically tuning the prompt, our proposed system is able to adequately adapt the final sculpture it creates.
@@ V Conclusion
In this work, we presented the first fully autonomous coarse-to fine and text-to-3D clay sculpting method. While our framework is not yet able to match human performance for this complex text-to-3D task, it is able to create a range of visually identifiable shapes, and even enable semantically tuning the final sculpture output. The primary limitation of the shapes created by our proposed system is that of surface roughness and imperfections that remain after grasp-based refinement actions. Through our results and human evaluation analysis, we believe that two key areas to explore for future work are that of more complex robot end-effectors to enable better finishing strategies (such as smoothing the clay surface) and to develop more reliable quantitative metrics to evaluate the quality of the final shapes created by these systems.
VI Acknowledgements

### v2

@@ IV-B Analysis of CLIP Scores for Clay Sculpting
Figure 4: Scatter plot with line of best fit for the CLIP and PointCLIP-v2 cosine similarity of text and image/point cloud embeddings of 10 human trajectories creating each shape in clay. The line of best fit’s slope for each shape and prompt shows how well the CLIP or PointCLIP-v2 score correlates with our human oracle-created shapes and varying prompts.
@@ IV-C Human Survey
TABLE II: Full Pipeline Evaluations. The Image and text CLIP scores are between 0 and 1. The human evaluations (denoted with H.) and LLM ratings are between 1 and 5.
@@
CLIP $\uparrow$ LLM $\uparrow$ H. Prompt $\uparrow$ H. Quality $\uparrow$ CLIP $\uparrow$ LLM $\uparrow$ H. Prompt $\uparrow$ H. Quality $\uparrow$
X 0.278 $\pm$ 0.014 2.33 $\pm$ 0.47 4.63 $\pm$ 0.70 3.75 $\pm$ 0.97 0.295 $\pm$ 0.008 2.67 $\pm$ 0.47 4.15 $\pm$ 0.53 3.69 $\pm$ 0.82
line 0.271 $\pm$ 0.009 2.67 $\pm$ 0.47 4.38 $\pm$ 0.70 3.38 $\pm$ 0.99 0.289 $\pm$ 0.002 3.33 $\pm$ 0.47 4.62 $\pm$ 0.63 4.15 $\pm$ 0.77
flower 0.294 $\pm$ 0.004 3.00 $\pm$ 0.00 3.13 $\pm$ 1.05 2.88 $\pm$ 0.78 0.302 $\pm$ 0.021 3.00 $\pm$ 0.00 4.31 $\pm$ 0.61 4.00 $\pm$ 1.11
column 0.270 $\pm$ 0.003 3.33 $\pm$ 0.47 2.88 $\pm$ 1.17 2.63 $\pm$ 1.22 0.306 $\pm$ 0.005 2.33 $\pm$ 0.47 3.31 $\pm$ 1.14 3.00 $\pm$ 1.04
pyramid 0.275 $\pm$ 0.009 3.00 $\pm$ 0.00 1.62 $\pm$ 0.74 1.54 $\pm$ 0.93 0.289 $\pm$ 0.007 3.00 $\pm$ 0.00 4.75 $\pm$ 0.42 3.75 $\pm$ 0.66
airplane 0.276 $\pm$ 0.009 3.00 $\pm$ 0.00 2.31 $\pm$ 0.82 1.85 $\pm$ 0.77 0.306 $\pm$ 0.005 2.67 $\pm$ 0.47 3.88 $\pm$ 0.93 3.13 $\pm$ 1.27
chair 0.297 $\pm$ 0.013 2.00 $\pm$ 0.00 1.85 $\pm$ 0.77 1.50 $\pm$ 0.76 0.327 $\pm$ 0.007 2.67 $\pm$ 0.47 3.00 $\pm$ 0.71 2.75 $\pm$ 0.97
pottery 0.275 $\pm$ 0.018 3.00 $\pm$ 0.00 1.39 $\pm$ 0.63 1.39 $\pm$ 0.63 0.297 $\pm$ 0.002 3.00 $\pm$ 0.00 4.02 $\pm$ 0.95 3.00 $\pm$ 1.23
all shapes 0.279 $\pm$ 0.014 2.79 $\pm$ 0.50 2.54 $\pm$ 1.37 2.18 $\pm$ 1.20 0.301 $\pm$ 0.015 2.83 $\pm$ 0.47 4.02 $\pm$ 0.95 3.50 $\pm$ 1.11
Figure 5: The human oracle is required to follow the same process of coarse-to-fine sculpting using their hands. The choice of camera orientation for each shape was to best visualize the full sculpture (i.e. top-down versus isometric viewpoint).
@@ IV-D Text-to-3D Shape Goals
To fully evaluate our proposed system, we conducted three real-world experiments for each shape prompt for the ’X’, ’line’, ’flower’, ’column’, ’pyramid’, ’airplane’, ’chair’, and ’pottery’ shape goals. A visualization of the final shapes created by our system is shown in Figure 5. As we present the first text-to-3D real-world shaping system, there are no existing baselines to compare to, thus we focus on comparisons to a human oracle. The quantitative results of the final CLIP cosine similarity, an LLM shape rating prompt using the same point cloud-based representation as our sub-goal module, and human evaluations for our system compared to a human oracle are shown in Table II. The shapes created by our pipeline are often visually similar to that of the human oracle, but with some imperfections such as indentations from the gripper. We can see from the image-based CLIP scores, the human oracle creates higher quality shapes than our proposed method (Welch’s t-test, $p=7.37e^{-6}$, $t=5.05$), though in absolute CLIP value our proposed method performs comparably for the ’X’, ’line’, and ’flower’ shapes. Similarly, the LLM consistently rated the human oracle point clouds higher than the point clouds of our method (Welch’s t-test, $p=0.77$, $t=0.29$). These patterns hold for human evaluation scores as well, with the human oracle being rated higher in both prompt similarity (Welch’s t-test, $p=1.47e^{-13}$, $t=8.14$) and shape quality (Welch’s t-test, $p=1.04e^{-11}$, $t=7.32$). However, we did find that human evaluators rated the ’X’ higher than the human oracle ’X’, though generally the evaluators favored the human oracle-created shapes. While there is a clear statistical significance between the human oracle and full pipeline performance on both human rating metrics, the human oracle itself did not receive a near-perfect score. The survey respondents struggled to quantify the shapes ’column’ and ’chair’, and to a lesser extent ’airplane’ and ’pottery’. It is very difficult to quantify the concept of how much a shape is a particular shape. To better investigate this difficulty in human evaluator consistency, we conducted a Krippendorff analysis [44] of the inter-annotator agreement across all shapes created by both the human oracle and our pipeline. For the ordinal data of our individual Likert-type questions, the alpha scores for the shape quality and prompt similarity were $\alpha_{quality}=0.533$, and $\alpha_{prompt}=0.544$, respectively. For the Krippendorff analysis, a score of 0 signifies that the level of agreement is consistent with randomly answering the questions, while a score of 1 means the reviewers perfectly agreed. The scores for both of our questions were approximately 0.5, which demonstrates that the responses are not random, but the ratings have substantial noise and the reviewers are not able to fully agree. This further emphasizes the necessity to explore better strategies to quantify performance for the clay shaping task in future work, as even human evaluations have disagreement about human-created shapes.
@@ IV-E Semantically Tuning Shape Goals
Figure 6: By semantically tuning the prompt, our proposed system is able to adapt the final sculpture it creates.
@@ V Conclusion
In this work, we present the first fully autonomous coarse-to fine and text-to-3D clay sculpting method. While our framework is not yet able to match human performance for this complex text-to-3D task, it is able to create a range of visually identifiable shapes, and even enable semantically tuning the final sculpture output. The primary limitation of the shapes created by our proposed system is that of surface roughness and some imperfections that remain after refinement actions. Additionally, the placement strategy for the individual chunks of clay limits the variety of final shapes our system is able to create, as there must always be supporting pieces below. Through our results and human evaluation analysis, we believe that the key areas to explore for future work are to build more complex robot end-effectors to enable better finishing strategies (such as smoothing the clay surface) and to develop more reliable quantitative metrics to evaluate the quality of the final shapes created by these systems without human evaluations.
VI Acknowledgments